If the perimeter of a rectangle is 100 meters, and the lengh is 1.5 times the width, find the area of the rectangle in square meters. Round to two decimal places.
Answer by srinivas.g(540)

Perimeter =100 meters
Let width be w meters
Then length (l)= 1.5*w
Perimeter =2(l+w)
100=2(1.5 w+w)
100=2(2.5w)
100=5 w
Divide with 5 on both sides
100/5=5w/5
20=w
Width =20 meters
Length=1.5×20
=30 meters
Area=length×width
=30×20
=600 square meters